This Algebra is making me crazy. When solving by elimination method, I get mixed up, for example:
5x + 2y = -13
7x – 3y = 19
Do I find the common denominator for y? I would multiply the top equation by 3 and the bottom equation by 2.
3(5x + 2y) = 3 (-13)
2(7x – 3y) = 2 (19)
15x + 6y = -39
14x – 6y = 38
Do I add or subtract?
29x = -1 0r 1x =-1
x = -1/29 or x -1/1 or just -1
I know that whatever I get for x, I then replace the x in either of the first two equations to solve for y. I would then have the ordered pair for the solution to the system.
Thank you so much.

Answer by stanbon(75887)

When solving by elimination method, I get mixed up, for example:
5x + 2y = -13
7x – 3y = 19
Do I find the common denominator for y? I would multiply the top equation by 3 and the bottom equation by 2.
3(5x + 2y) = 3 (-13)
2(7x – 3y) = 2 (19)
-------
Good at this point.
---------
15x + 6y = -39
14x – 6y = 38
-------
Good at this point.
-------------
Do I add or subtract?
----
You add so the "y-terms" will be eliminated:
29x = -1 0r 1x =-1
x = -1/29 or x -1/1 or just -1
I know that whatever I get for x, I then replace the x in either of the first two equations to solve for y. I would then have the ordered pair for the solution to the system.
---------
Correct
-------
Cheers,
Stan H.